Matome borrows R4 050 for eight months from a lender who charges a 11% discount rate. How much money does Matome receive?


1
Expert's answer
2021-05-18T13:30:46-0400

"FV=PV(1+i)^n=4050(1+0.11)^8=R\ 9333.38"